OpenStudy (anonymous):

Could someone PLEASE help me?! Match each system of linear equations with the quadrant in which the solution lies. Match Term Definition 3x – 2y = 10 5x + 3y = 4 A) III ---------- 3x + 2y = 14 x – 3y = –10 B) I ------------- x + 2y = 8 4x + 6y =22 C) II -------------- 2x + 3y = –12 4x – 2y = –2 D) IV

OpenStudy (displayerror):

I would assume the following procedure would apply to the remaining three: For the linear equations \[3x-2y=10\] and \[5x+3y=4\] Put the two equations in terms of y=, so \[3x-2y=10 \rightarrow 2y=3x-10 \rightarrow y= \frac{3x-10}{2}\] \[5x+3y=4 \rightarrow 3y=4-5x \rightarrow y=\frac{4-5x}{3}\] Because both are of the form "y=", we can set the two equations equal to one another: \[y= \frac{3x-10}{2} = \frac{4-5x}{3}\] Solve for x by performing cross multiplication: \[3*(3x-10) = 2*(4-5x)\] \[9x-30 = 8-10x\] \[19x = 38 \rightarrow x=\frac{38}{19}\] Now plug that value of x into your original equation and find the y-value corresponding to that x that we just found. After that, you should have a coordinate (x,y), and looking at that, you should know which quadrant the two lines intersect.